OpenStudy (anonymous):

If a coin is flipped 10 times, what is the probability that it will show all heads or all tails?

OpenStudy (ikram002p):

p(h)^10+p(t)^10 (0.5)^10+(0.5)^10=2(0.5)^10

OpenStudy (anonymous):

a: 1/2 b: 1/512 c: 1/1,022 d: 1/1,024 e: 1/2,048

OpenStudy (ikram002p):

its as i said p(h)^10+p(t)^10 (1/2)^10+(1/2)^10=2(1/2)^10=1/2,048